OpenStudy (anonymous):

Find the following expression by multiplying. (a + b)^3

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

might should memorize this, it is not that long\[(a+b)^3=a^3+3a^2b+3ab^2+b^3\]

OpenStudy (anonymous):

@misty1212 how would i do the foil method with this?

OpenStudy (misty1212):

there is no such thing as foil there is multiply

OpenStudy (misty1212):

\[(a+b)(a+b)(a+b)=(a^2+2ab+b^2)(a+b)=a^3+3a^2b+3ab^2+b^3\]

OpenStudy (anonymous):

Thank you

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]